OpenStudy (anonymous):

How do I find the integral of 1/(sqrt(1+sqrt(x))) dx

ganeshie8 (ganeshie8):

Have you tried \(\large 1+\sqrt{x} = u^2\) ?

OpenStudy (anonymous):

Ah I didnt think about to square u, let me try.

OpenStudy (anonymous):

Uuhm, so \[du ^{2}=\frac{ dx }{ 2\sqrt{x} }\]

OpenStudy (anonymous):

We didnt learn about squaring u's before so I'm not sure how it works excactly

OpenStudy (anonymous):

no its 2du=dx/2x^1/2

OpenStudy (anonymous):

Ok so I got \[2du=\frac{ dx }{ 2\sqrt{x} }\] How do I get \[\sqrt{1+\sqrt{x}}\] into \[\frac{ dx }{ 2\sqrt{x} }\]

OpenStudy (anonymous):

I mean \[\frac{ dx }{ \sqrt{1+\sqrt{x}} }\]

OpenStudy (anonymous):

\[u got the answer\]

OpenStudy (anonymous):

I don't see how im there yet :P I need to get \[\frac{ dx }{ \sqrt{1+\sqrt{x}} }\] into \[\frac{ dx }{ 2\sqrt{x} }\]

ganeshie8 (ganeshie8):

\[\large 1+\sqrt{x} = u^2 \implies \dfrac{1}{2\sqrt{x}}dx = 2u du \implies dx = 4u(u^2-1)du\]

ganeshie8 (ganeshie8):

the integral becomes : \[\large \int \dfrac{4u(u^2-1)du}{\sqrt{u^2}}\]

OpenStudy (anonymous):

How did you got from \[\frac{ 1 }{ 2\sqrt{x} }dx =2udu\] to \[dx =4u(u^2-1)du\]

ganeshie8 (ganeshie8):

substitute \(\large \sqrt{x} = u^2-1\)

OpenStudy (anonymous):

Ohyea ofcourse :P Thanks alot!!

OpenStudy (anonymous):

Sorry but this one is really hard. Further I got the integral and is \[\frac{ 4 }{ 3 }u ^{3}-4x\] which becomes \[\frac{ 4 }{ 3 }(1+\sqrt{x})^{\frac{ 3 }{ 2 }}-4x\] Because \[u ^{2}=1+\sqrt{x}\] so \[u ^{3}=(1+\sqrt{x})^{\frac{ 3 }{ 2 }}\] I must have done something wrongg but i dont knwo what